User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!

Meeting 7th December

Chair: Tiege

Present: Jonny, Patrick, Ross, Arwin, Mas, Liz, Ellen, Lucia, Jamie

Apologies: Jesse, Bryn

Absent: Lucas.

General Issues:

  • 26m Chaos
    • Lightening caused some issues, knocked out serial communications to several systems. Spurious alarms (on-going), live page is not working correctly (Should be fixed), X-axis cryo pipe failed (Fixed now), lacking wind-sensor (on-going), however, auto wind-stow script should still be working.
  • Cont17 progress, thoughts or concerns?
    • S/X receiver pressure had issues, potentially was linked to the cryo leak, should be fixed.
  • Bedroom (Lounge), heater
    • If you use it, turn it off!!! Don't burn down the building.
  • Atifur start time
    • Sometime in the new year.
  • Internet outage for Sunday
    • This is going ahead, internet will drop for some duration, hopefully less than an hour. Will be affecting MtP also. Jamie will try and get them to give us more detailed plans.

Issues in Handover Notes:

  • Semi-regular “ previous source in this schedule not reached before new source was commanded.” errors on Hobart 26m during Cont17.
    • Decelleration issues, Resistor died, was replaced with another one. Majority of the time not an issue, but drives are struggling keeping up with this scheduler.
  • Hobart 26m not stowing correctly during wind stow.
    • Should be fixed, keep an eye on it. Jamie will look into it more.
  • Keep an eye on receiver pressure, needed to go to MtP for CRDS92 and turbo pump it.
    • Related to cryo issues
  • Yg Anemometer
    • Power cycling using the ipswitch for the anemometer and restarting Monica will fix this.
/home/www/auscope/opswiki/data/pages/operations/agenda2017_12_07.txt · Last modified: 2017/12/19 17:38 by Patrick Yates